# Measuring Iterative Temporal Reasoning with [Uncaptioned image]Time Puzzles

### v2

@@ A.2 Algorithm Implementation Details
As mentioned in Section 2.2, we generate each puzzle by sampling a hidden seed date and selecting $N$ natural-language constraints whose conjunction yields a user-defined number of solutions. To ensure diversity and computational efficiency, our pipeline utilizes a comprehensive taxonomy of temporal facts and sorts constraints by Information Gain (IG) during verification.
@@ 6 $\mathcal{A}\leftarrow\mathcal{A}\cap\mathcal{C}(t)$;
Appendix B Evaluation Metrics
In the paper, we primarily use exact match accuracy to measure model performance to emphasize our focus on precise iterative reasoning, but we also consider F1 and Jaccard Index to enhance our evaluation. We describe their computation and the empirical difference between exact match accuracy and F1/Jaccard Index.
B.1 Metric Computation
Let $Y_{i}$ denote the gold solution set and $\hat{Y}_{i}$ the predicted set for instance $i$.
Exact Match Accuracy
$\mathrm{EM}_{i}=\mathbb{I}\left[Y_{i}=\hat{Y}_{i}\right]$ (3)
Precision, Recall, and F1
$\mathrm{Precision}_{i}=\frac{|Y_{i}\cap\hat{Y}_{i}|}{|\hat{Y}_{i}|},\quad\mathrm{Recall}_{i}=\frac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}|},$ (4)
$\mathrm{F1}_{i}=\frac{2\,\mathrm{Precision}_{i}\,\mathrm{Recall}_{i}}{\mathrm{Precision}_{i}+\mathrm{Recall}_{i}}.$ (5)
By convention, when $Y_{i}=\hat{Y}_{i}=\emptyset$, all three metrics are set to 1; when exactly one of the sets is empty, all three metrics are set to 0.
Jaccard Index
$\mathrm{Jaccard}_{i}=\begin{cases}1&\text{if }Y_{i}=\hat{Y}_{i}=\emptyset\\ \dfrac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}\cup\hat{Y}_{i}|}&\text{otherwise}\end{cases}$ (6)
Final reported scores are computed via macro-averaging:
$\mathrm{Metric}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{Metric}_{i},$ (7)
where $N$ is the number of evaluation instances.
Mean Std Dev % Equal Spearman
EM Vs F1 -7.07 20.52 87.98 0.92
EM Vs JI -5.40 16.41 87.98 0.92
Table 3: Exact match (EM) accuracy versus F1/Jaccard Index (JI) based on all the experiments ran on this study (over 19,000 data points).
B.2 Exact Match Versus F1/Jaccard Index
Table 3 compares exact match (EM) accuracy with F1 and Jaccard Index (JI). The mean absolute differences between EM and F1 and between EM and JI are 7.07 and 5.40, respectively. The relatively large standard deviations are expected, as EM is binary at the instance level. Nevertheless, in 87.98% of cases, all three metrics agree, indicating that models typically either solve a puzzle completely or fail entirely. Consistent with this observation, Spearman correlations between EM and the other metrics are high ($r=0.92$).
@@ Appendix C Models
Providing Web Search and Code Interpreter simultaneously does not guarantee that both tools are invoked during generation, even when tool choice is set as required. To control costs and ensure consistent retrieval, we therefore reuse the web-search context collected from the Web Search–only setting and prepend it to the prompt for GPT models, while enabling Code Interpreter with tool choice set as required.
Appendix D Additional Results
Figure 4 reports average exact match accuracy across solution counts for the remaining models (for which web search is not implemented) under both implicit and explicit constraints.
Figure 4: Average exact-match accuracy across solution counts for other models evaluated without web search.
Appendix E Prompts
We use the following prompt template. The placeholder {constraints} is replaced with the instance-specific constraint set at inference time.

### v3

@@ A.2 Algorithm Implementation Details
We generate each puzzle by sampling a hidden seed date and selecting $N$ natural-language constraints whose conjunction yields a user-defined number of solutions. To ensure diversity and computational efficiency, our pipeline utilizes a comprehensive taxonomy of temporal facts and sorts constraints by Information Gain (IG) during verification.
@@ Appendix B Models
Providing Web Search and Code Interpreter simultaneously does not guarantee that both tools are invoked during generation, even when tool choice is set as required. To control costs and ensure consistent retrieval, we therefore reuse the web-search context collected from the Web Search-only setting and prepend it to the prompt for GPT models, while enabling Code Interpreter with tool choice set as required.
Appendix C Prompts
C.1 Default Prompt
We use the following prompt template in the main experiments in Section 3. The placeholder {constraints} is replaced with the instance-specific constraint set at inference time.
@@ C.1 Default Prompt
C.2 Prompt with Step-by-step Reasoning Instruction
The following prompt is used in the follow-up experiment in Section 4, which builds on the default prompt by inserting detailed step-by-step reasoning instruction.
From the time-related constraints below, determine all valid date(s) (if any) that satisfy them.
Depending on the conditions, the result may be zero, one, or multiple dates.
Unless otherwise stated, interpret all constraints using the Gregorian calendar.
Carefully follow these step-by-step reasoning instructions:
1) Read and restate the goal
- You must find ALL Gregorian calendar date(s) that satisfy EVERY constraint.
- The answer can be none, one, or many dates.
2) Normalize definitions you must use
- Seasons:
* Winter = Dec, Jan, Feb
* Spring = Mar, Apr, May
* Summer = Jun, Jul, Aug
* Autumn = Sep, Oct, Nov
- If constraints mention weekday, leap year, end of month, etc., use standard Gregorian rules.
3) Extract constraints into a checklist
- Parse the text and list each constraint as a separate item.
- If any constraint contains an OR, split into separate branches.
- If constraints contain ranges (e.g., "between 2010 and 2015"), record inclusivity/exclusivity exactly as stated.
4) Determine the search space (candidate window)
- Identify the tightest restriction on:
* Year(s) (specific year, range, century, decade)
* Month(s) (specific month(s), season, quarter)
* Day(s) (day number, end of month, first Monday, etc.)
- Use these to form an initial set of candidate dates or a candidate-generating rule.
5) Convert each constraint into a filter function
For each constraint, translate it into a precise test such as:
- Year filter (e.g., year = 2024; year in [1990, 1999])
- Month/season filter (e.g., month in {{6,7,8}})
- Day-of-month filter (e.g., day = 1; day >= 28)
- Weekday filter (e.g., weekday = Monday)
- Relative-date relations (e.g., "the day after X", "two weeks before Y")
- Calendar properties:
* Leap year rule: leap if divisible by 4, except centuries not divisible by 400.
* Month lengths: 30/31 days; February 28 or 29 depending on leap year.
6) Generate candidates in the most restrictive way
- Prefer generating candidates from the most specific constraints first:
* If a specific date is implied, compute it directly.
* If a specific year + season is given, iterate only those months.
* If a weekday pattern exists ("first Monday of May"), generate those directly instead of brute-forcing all days.
7) Apply constraints systematically
- For each candidate date, check constraints one by one.
- Eliminate candidates immediately when one constraint fails.
- Keep a running list of remaining candidates.
8) Handle OR and branching logic correctly
- If constraints are like: (A OR B) AND C:
* Compute candidates satisfying A and C
* Compute candidates satisfying B and C
* Take the union of those results (deduplicate).
9) Check for hidden contradictions early
- Before enumerating too much, look for impossibilities, e.g.:
* Month is February but day is 30
* Season says Winter but month says July
* Date required to be both Monday and Tuesday
* Before 2000 AND after 2010
- If a contradiction is found, conclude no solution.
10) Ensure completeness (dont miss valid dates)
- Confirm you explored all years/months allowed by the constraints.
- If you used branching, ensure all branches were evaluated.
11) Sort and format final answers
- Sort valid dates chronologically.
- Output dates in YYYY-$M$-$D$ format.
- If multiple, separate by commas and a space.
The constraints are as follows:
{constraints}
After thorough consideration, end your response on a new line with "MY ANSWER: "
followed by the valid date(s) in the format "YYYY-$M$-DD".
If there are multiple valid dates, list them separated by commas.
If no valid date exists, respond with "MY ANSWER: None".